# SplatSSC: Decoupled Depth-Guided Gaussian Splatting for Semantic Scene Completion

### v1

@@ Introduction
To evaluate the effectiveness of our SplatSSC, we conduct extensive experiments on the high-quality indoor datasets Occ-ScanNet and Occ-ScanNet-mini (Yu et al. 2024a). Details about datasets, implementation, and evaluation metrics are included in our supplementary material.
@@
The main results on the Occ-ScanNet and Occ-ScanNet-mini benchmarks are summarized in Table 1. Our SplatSSC achieves SOTA performance, demonstrating strong robustness and fine-grained scene understanding on both benchmarks. For Occ-ScanNet, SplatSSC achieves 62.83% IoU and 51.83% mIoU, surpassing the previous SOTA RoboOcc (Zhang et al. 2025) by a substantial margin of 6.35% and 4.16%, respectively. The per-class analysis further highlights the consistent improvements brought by SplatSCC across diverse categories, from large structural elements (e.g., walls and floor) to fine-grained objects (e.g., sofas and chairs). These results underscore the strength of our synergistic design. The depth-guided initialization facilitates accurate geometric reconstruction, while our DGA ensures sharp semantic boundaries. As illustrated in the qualitative examples in Figure 5, SplatSSC yields superior 3D scene perception capabilities that surpass the previous paradigm.
Figure 5: Qualitative results on the Occ-ScanNet-mini dataset. Our method achieves superior performance in scene completion and target object recall compared to the existing approaches.
@@
[…]
@@
[…]
@@
The contribution of our GMF module is further validated in Table 5. The results highlight a dramatic impact of GMF on refining the geometric prior. When applied to a frozen Depth-Anything-V2 (DAv2) backbone, our GMF module boosts the $\delta_{1}$ score by a remarkable 0.906. Furthermore, GMF demonstrates its capability to enhance even a strong and fine-tuned Depth-Anything-V2 (FT-DAv2) (Wu et al. 2025), further pushing the $\delta_{1}$ score to a new best of 0.993. This confirms that our GMF is a powerful and versatile feature refiner, essential for generating high-quality geometric representations.
Efficiency Analysis
[…]
@@ Conclusion
Despite its outstanding performance, we acknowledge several limitations that offer avenues for future work, which will be discussed in the supplementary material.
@@
Occ-ScanNet (Yu et al. 2024a) comprises 45,755 training frames and 19,764 validation frames, annotated with 12 semantic classes, with one representing free space and eleven corresponding to specific categories, including ceiling, floor, wall, window, chair, bed, sofa, table, television, furniture, and generic objects. The ground truth is provided as a voxel grid covering a 4.8m×$\times$4.8m×$\times$2.88m region in front of the camera, discretized into a resolution of 60$\times$60$\times$36. This dataset serves as the benchmark for training and evaluating local occupancy prediction. A smaller variant, Occ-ScanNet-mini, is also available, containing 5,504 training and 2,376 validation frames.
@@
[…]
@@
In our framework, the image encoder employs a pretrained EfficientNet-B7 (Tan and Le 2019) as backbone, while the depth branch utilizes a frozen fine-tuned Depth-Anything-V2 (Wu et al. 2025) model. For both training stages, we use the AdamW optimizer (Loshchilov and Hutter 2019) with a weight decay of 0.01. We apply a learning rate multiplier of 0.1 to the backbone. All input images are processed at a resolution of 480×640480$\times$ 640480 × 640.
@@
In the first stage, we exclusively pretrain our depth branch to establish a robust geometric prior. The down-sampled grid for our GMF has a shape of 30×4030$\times$ 4030 × 40. We employ a cosine learning rate schedule with a 1000-iteration warmup, setting the peak learning rate to $6\times 10^{-4}$. The model is trained for 10 epochs on the Occ-ScanNet dataset using 2 NVIDIA 3090 GPUs with a per-GPU batch size of 2 (total batch size of 4).
@@
In the second stage, we train the full SplatSSC model, initializing the depth branch with weights from stage one. The 30×4030$\times$ 4030 × 40 down-sampled grid generates an initial set of 1200 Gaussian primitives, with their scales initialized in the range [0.01m, 0.16m]. We train the model on 4 NVIDIA 4090 GPUs with a per-GPU batch size of 2, resulting in a total batch size (bs) of 8. The learning rate follows a cosine schedule with a 1000-iteration warmup, and the peak learning rate is determined by a linear scaling rule: $2\times 10^{-4}\cdot(\text{bs}/2)$. The model is trained for 10 epochs on the full Occ-ScanNet dataset and for 20 epochs on the Occ-ScanNet-mini subset.
@@
In Figure 8, we present additional qualitative results for 3D occupancy prediction on the Occ-ScanNet validation set. These examples further demonstrate SplatSSC’s ability to reconstruct detailed object shapes within diverse and complex scenes accurately.
Figure 8: Further visualization on the Occ-ScanNet dataset.
Furthermore, Figure 7 provides more visualizations on the Occ-ScanNet-mini validation set. Alongside the final occupancy predictions, we also include per-frame visualizations of the intermediate 3D semantic Gaussians. These visualizations offer insight into our model’s underlying sparse, object-centric scene representation, illustrating how it efficiently models the scene’s geometry and semantics before rendering the final dense grid.
@@ Discussion
This experiment validates our finding that SplatSSC’s performance is subject to a distinct threshold regarding its training hyperparameters, with results shown in Table 7. This effect is visible when comparing performance at different batch sizes. At a total batch size of 2, our model’s performance is substantially limited to 36.09% mIoU. However, upon increasing the batch size to 4, the mIoU jumps dramatically to 45.32%, reaching a competitive level. This demonstrates that a batch size of at least 4 is necessary for effective optimization. Beyond this threshold, performance continues to scale robustly, with the best results achieved at a batch size of 8. In contrast, the baseline EmbodiedOcc (Wu et al. 2025) exhibits only modest and linear gains. It is also worth noting that EmbodiedOcc was designed for a per-GPU batch size of one, making extensive scaling less applicable. This highlights that the observed threshold effect is a unique characteristic of our model’s interactive primitive optimization.
@@
The current SplatSSC framework is designed to operate on a per-frame basis, excelling at generating a high-quality scene representation from a single view. However, this design presents a scalability challenge when extending to global scene perception. A naive extension of simply accumulating primitives from consecutive frames would lead to an unbounded growth in their total number, causing a rapid escalation in both memory and computational costs. This limitation reveals a critical need for a scalable online primitive management strategy that leverages both pruning and fusion techniques to prevent unbounded growth in memory and computation. We leave this as a promising direction for future work and will validate it on global scene benchmarks (Wu et al. 2025; Wang et al. 2024b).
@@
A primary direction is adapting SplatSSC for large-scale outdoor environments, particularly for applications like autonomous driving. Unlike methods that rely on dense grids (Wei et al. 2023; Zhang, Zhu, and Du 2023) or random initialization across a predefined volume (Huang et al. 2025), our depth-guided approach naturally focuses computation on observed surfaces. This inherent efficiency makes it exceptionally well-suited for sparse and large-scale settings. To fully realize this potential, the fixed volumetric grid could be replaced with more flexible spatial data structures, such as hash-encoded grids (Deng et al. 2025b), to support unbounded scenes. This extension would also need to address the challenges unique to this domain, such as managing a dynamically growing set of primitives and handling the presence of dynamic objects.
@@
Moving beyond passive perception, a critical frontier in 3D vision is to build representations that support active interaction, a central theme in embodied and spatial intelligence (Wang et al. 2024b; Halacheva et al. 2025). Applying SplatSSC in embodied AI requires moving from single-frame perception to building a persistent and interactive world model. This demands a higher level of detail than is currently captured; for instance, an agent needs not just a semantic label for a “door”, but also precise geometric information about its handle for manipulation. This may necessitate using a larger number of Gaussians or a finer-grained semantic taxonomy. Furthermore, it requires a robust online framework where the agent can continuously fuse new observations (Deng et al. 2025a), prune outdated information, and refine its Gaussian-based world map in real-time.
@@
The source code and trained models associated with this paper will be made publicly available upon acceptance. The repository will be hosted on GitHub, and the specific URL will be provided here.

### v2

@@ Introduction
To evaluate the effectiveness of our SplatSSC, we conduct extensive experiments on the high-quality indoor datasets Occ-ScanNet and Occ-ScanNet-mini (yu2024monocular). Details about datasets, implementation, and evaluation metrics are included in our supplementary material from our code.
@@
The main results on the Occ-ScanNet and Occ-ScanNet-mini benchmarks are summarized in Table 1. Our SplatSSC achieves SOTA performance, demonstrating strong robustness and fine-grained scene understanding on both benchmarks. For Occ-ScanNet, SplatSSC achieves 62.83% IoU and 51.83% mIoU, surpassing the previous SOTA RoboOcc (zhang2025roboocc) by a substantial margin of 6.35% and 4.16%, respectively. The per-class analysis further highlights the consistent improvements brought by SplatSSC across diverse categories, from large structural elements (e.g., walls and floor) to fine-grained objects (e.g., sofas and chairs). These results underscore the strength of our synergistic design. The depth-guided initialization facilitates accurate geometric reconstruction, while our DGA ensures sharp semantic boundaries. As illustrated in the qualitative examples in Figure 5, SplatSSC yields superior 3D scene perception capabilities that surpass the previous paradigm.
Figure 5: Qualitative results on the Occ-ScanNet-mini dataset. Our method achieves superior performance in scene completion and target object recall compared to others.
@@
We analyze the impact of primitive count and scale range in Table 2, revealing a clear trade-off between performance and efficiency. Our setting achieves the highest semantic accuracy of $48.87\%$ mIoU with a remarkably compact configuration of just 1200 primitives. Increasing the count to 4800 and 19200 yields marginal gains in geometric completeness but incurs higher latency and lower mIoU. The choice of scale range is equally critical. Excessively large ranges degrade accuracy and trigger Out-of-Memory (OOM) failures under dense configurations, likely due to overlaps among oversized primitives. In contrast, a moderate range [0.01, 0.16] offers the best trade-off, effectively capturing both global layouts and fine-grained details with minimal redundancy.
@@
We evaluate the impact of our key components, GMF and DGA, in Table 3. The analysis first highlights the necessity of a tailored aggregation mechanism. The standard GF.agg (huang2024gaussianformer) nearly fails in our sparse setting, yielding a prohibitively low 10.45% mIoU. While the more advanced GF2.agg (huang2025gaussianformer2) performs significantly better, our DGA still surpasses it by over $2.8\%$ in both IoU and mIoU. This confirms that “floaters” are the key bottleneck in sparse splatting, and DGA is crucial for efficient and robust aggregation. The proposed GMF is equally important, as replacing it with a naive depth-aware baseline (wu2024embodiedocc) built on Depth-Anything-V2 causes a substantial drop by more than $11\%$ in both geometries and semantics, even when paired with our DGA. The degradation becomes more severe with other aggregators, leading to a near-collapse in performance. This demonstrates the necessity of structured geometric priors for generating informative primitives.
@@
The contribution of our GMF module is validated in Table 5. The results highlight a dramatic impact of GMF on refining the geometric prior. When applied to a frozen Depth-Anything-V2 (DAv2) backbone, GMF boosts the $\delta_{1}$ score by a remarkable 0.906. Furthermore, it demonstrates its capability to enhance a fine-tuned Depth-Anything-V2 (FT-DAv2) (wu2024embodiedocc), pushing the $\delta_{1}$ score to a new best of 0.993. This confirms that our GMF is a powerful and versatile feature refiner, essential for generating high-quality geometric representations.
Efficiency Analysis.
Beyond accuracy, we evaluate the computational efficiency of SplatSSC against EmbodiedOcc, with results detailed in Figure 6. Our method demonstrates superior efficiency despite a negligible $0.19\%$ increase in parameter count. Specifically, SplatSSC achieves a $9.32\%$ reduction in inference latency and a $9.64\%$ decrease in memory usage. This advantage is primarily attributed to our sparse design, which operates on significantly fewer primitives than prior works.
@@ Conclusion
Despite its outstanding performance, we acknowledge several limitations that offer avenues for future work as discussed in the supplementary material from our code.
Acknowledgments
This work was supported by National Research Foundation of Singapore Medium-sized Centre for Advanced Robotics Technology Innovation and Ministry of Education, Singapore, under AcRF TIER 1 Grant RG64/23.
@@ Acknowledgments
  • We provide further proof and analysis to support the necessity of our two-stage training.
@@
Occ-ScanNet (yu2024monocular) comprises 45,755 training frames and 19,764 validation frames, annotated with 12 semantic classes, with one representing free space and eleven corresponding to specific categories, including ceiling, floor, wall, window, chair, bed, sofa, table, television, furniture, and generic objects. The ground truth is provided as a voxel grid covering a $\times4.8m$$\times$2.88m region in front of the camera, discretized into a resolution of $\times60$$\times$36. This dataset serves as the benchmark for training and evaluating local occupancy prediction. A smaller variant, Occ-ScanNet-mini, is also available, containing 5,504 training and 2,376 validation frames.
@@
Following common practice (cao2024monoscene; hu2024metric3d), we evaluate the final semantic scene completion performance using Intersection-over-Union (IoU) and mean IoU (mIoU). These metrics are computed exclusively within the current camera’s view frustum. To assess the quality of the geometric prediction in our depth branch, we employ three additional metrics: Chamfer $l_{1}$ distance (C-$l_{1}$), Root Mean Squared Error (RMSE), and accuracy under threshold ($\delta_{1}$(hu2024metric3d). For this geometric evaluation, the ground truth point cloud is generated by down-sampling the ground truth depth map using the indices from our GMF module, then projecting the valid depth points into the camera’s coordinate space.
@@
In our framework, the image encoder employs a pretrained EfficientNet-B7 (tan2019efficientnet) as backbone, while the depth branch utilizes a frozen fine-tuned Depth-Anything-V2 (wu2024embodiedocc) model. For both training stages, we use the AdamW optimizer (loshchilov2017decoupled) with a weight decay of 0.01. We apply a learning rate multiplier of 0.1 to the backbone. All input images are processed at a resolution of $480\times 640$.
@@
In the first stage, we exclusively pretrain our depth branch to establish a robust geometric prior. The down-sampled grid for our GMF has a shape of $30\times 40$. We employ a cosine learning rate schedule with a 1000-iteration warmup, setting the peak learning rate to $6\times 10^{-4}$. The model is trained for 10 epochs on the Occ-ScanNet dataset using 2 NVIDIA 3090 GPUs with a per-GPU batch size of 2 (total batch size of 4).
@@
In the second stage, we train the full SplatSSC model, initializing the depth branch with weights from stage one. The $30\times 40$ down-sampled grid generates an initial set of 1200 Gaussian primitives, with their scales initialized in the range [0.01m, 0.16m]. We train the model on 4 NVIDIA 4090 GPUs with a per-GPU batch size of 2, resulting in a total batch size (bs) of 8. The learning rate follows a cosine schedule with a 1000-iteration warmup, and the peak learning rate is determined by a linear scaling rule: $2\times 10^{-4}\cdot(\text{bs}/2)$. The model is trained for 10 epochs on the full Occ-ScanNet dataset and for 20 epochs on the Occ-ScanNet-mini subset.
@@
Figure 8: Qualitative visualization on the depth prediction.
Figure 9: Comparison of different training strategies.
@@
Further Visualization on SSC
In Figure 10, we present additional qualitative results on the Occ-ScanNet validation set. For each scene, we visualize both the intermediate 3D semantic Gaussians and the final dense occupancy prediction. The Gaussian views reveal that SplatSSC forms a sparse, object-centric representation that already captures well-aligned shapes and semantics for major structures (e.g., furniture and walls), while the rendered voxel grids show that this compact representation can be faithfully translated into detailed 3D completions.
Furthermore, Figure 7 provides more visualizations on the Occ-ScanNet-mini validation set using the same layout. Again, we show the per-frame 3D semantic Gaussians together with the final occupancy predictions. These examples highlight that our Gaussian scene representation remains stable and expressive in smaller and more cluttered environments, and that it consistently supports accurate reconstruction of fine-scale object geometry.
Further Visualization on Depth Branch
We provide additional qualitative visualizations of our depth branch in Figure 8. For each scene, we show the ground-truth depth at its naively downsampled $30\times 40$ version, followed by the output of our depth branch and FT-DAv2 at the same $30\times 40$ resolution, which can be used to initialize the initial Gaussian primitives in location.
Compared to FT-DAv2, our depth branch produces depth maps that are visibly closer to the ground truth. It better maintains a clearer separation between foreground objects and background surfaces.
These results indicate that our proposed depth branch can effectively refine the depth prior into a more structured and compact geometric representation, providing a stronger initialization for the subsequent Gaussian lifting stage.
Figure 10: Further visualization of predicted 3D Gaussians and voxels on the Occ-ScanNet dataset.
@@ Discussion
Analysis of Two-Stage Training Strategy
Figure 9 compares our two-stage schedule with an end-to-end training variant. In both cases, we use the SSC loss:
$\mathcal{L}_{\mathrm{ssc}}$ $=\mathcal{L}_{\mathrm{sem}}+\lambda_{1}\mathcal{L}_{\mathrm{scal}}^{\mathrm{prob}},$ (21)
$\mathcal{L}_{\mathrm{sem}}$ $=\lambda_{2}\mathcal{L}_{\mathrm{focal}}+\lambda_{3}\mathcal{L}_{\mathrm{lovasz}}.$ (22)
Our two-stage model in the second stage is optimized only with $\mathcal{L}_{\mathrm{ssc}}$, while the end-to-end baseline additionally uses the depth loss:
$\mathcal{L}_{\mathrm{d}}=\lambda_{4}\mathcal{L}_{\mathrm{huber}}^{\mathrm{depth}}+\lambda_{5}\mathcal{L}_{\mathrm{huber}}^{\mathrm{pts}}+\lambda_{6}\mathcal{L}_{\mathrm{grad}}.$ (23)
For a fair comparison, Figure 9 reports only $\mathcal{L}_{\mathrm{ssc}}$ together with the validation mIoU. In our experiments, we set $\lambda_{1}=\lambda_{6}=0.5$, $\lambda_{2}=100$, $\lambda_{3}=2$, $\lambda_{4}=20$, $\lambda_{5}=10$.
We observe that the end-to-end variant converges to a larger $\mathcal{L}_{\mathrm{ssc}}$ and a lower mIoU, whereas our two-stage schedule achieves a smaller loss and a higher mIoU. We attribute this gap to the absence of a precise depth prior in the end-to-end setting: jointly training the depth branch from scratch couples noisy depth supervision with SSC optimization and makes the Gaussian lifting less stable. In contrast, the first stage of our two-stage training learns a reliable depth prior under metric supervision, and the second stage then refines the Gaussian representation using $\mathcal{L}_{\mathrm{ssc}}$ only, which better aligns the pretrained depth cues with the SSC objective.
@@
This experiment validates our finding that SplatSSC’s performance is subject to a distinct threshold regarding its training hyperparameters, with results shown in Table 7. This effect is visible when comparing performance at different batch sizes. At a total batch size of 2, our model’s performance is substantially limited to 36.09% mIoU. However, upon increasing the batch size to 4, the mIoU jumps dramatically to 45.32%, reaching a competitive level. This demonstrates that a batch size of at least 4 is necessary for effective optimization. Beyond this threshold, performance continues to scale robustly, with the best results achieved at a batch size of 8. In contrast, the baseline EmbodiedOcc (wu2024embodiedocc) exhibits only modest and linear gains. It is also worth noting that EmbodiedOcc was designed for a per-GPU batch size of one, making extensive scaling less applicable. This highlights that the observed threshold effect is a unique characteristic of our model’s interactive primitive optimization.
@@
The current SplatSSC framework is designed to operate on a per-frame basis, excelling at generating a high-quality scene representation from a single view. However, this design presents a scalability challenge when extending to global scene perception. A naive extension of simply accumulating primitives from consecutive frames would lead to an unbounded growth in their total number, causing a rapid escalation in both memory and computational costs. This limitation reveals a critical need for a scalable online primitive management strategy that leverages both pruning and fusion techniques to prevent unbounded growth in memory and computation. We leave this as a promising direction for future work and will validate it on global scene benchmarks (wu2024embodiedocc; wang2024embodiedscan).
@@
A primary direction is adapting SplatSSC for large-scale outdoor environments, particularly for applications like autonomous driving. Unlike methods that rely on dense grids (wei2023surroundocc; zhang2023occformer) or random initialization across a predefined volume (huang2025gaussianformer2), our depth-guided approach naturally focuses computation on observed surfaces. This inherent efficiency makes it exceptionally well-suited for sparse and large-scale settings. To fully realize this potential, the fixed volumetric grid could be replaced with more flexible spatial data structures, such as hash-encoded grids (deng2025vpgs), to support unbounded scenes. This extension would also need to address the challenges unique to this domain, such as managing a dynamically growing set of primitives and handling the presence of dynamic objects.
@@
Moving beyond passive perception, a critical frontier in 3D vision is to build representations that support active interaction, a central theme in embodied and spatial intelligence (wang2024embodiedscan; halacheva2025gaussianvlm). Applying SplatSSC in embodied AI requires moving from single-frame perception to building a persistent and interactive world model. This demands a higher level of detail than is currently captured; for instance, an agent needs not just a semantic label for a “door”, but also precise geometric information about its handle for manipulation. This may necessitate using a larger number of Gaussians or a finer-grained semantic taxonomy. Furthermore, it requires a robust online framework where the agent can continuously fuse new observations (deng2025mne), prune outdated information, and refine its Gaussian-based world map in real-time.
@@
The source code and trained models associated with this paper will be released on GitHub upon acceptance, and the specific URL will be provided.